OpenStudy (anonymous):

convert 1/2*1% to a decimal

OpenStudy (anonymous):

did you try: 1/2 * 1/100 =

OpenStudy (anonymous):

\[\frac{1}{2} \times \frac{1}{100} = \frac{1}{200} \approx 0.005\] when you face (number) %<-- number/100 ^_^

OpenStudy (anonymous):

k ill try one

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

16/25= .0064

OpenStudy (anonymous):

watch out, it's 0.64 :)

OpenStudy (anonymous):

thanks